What is the acceleration of a 45kg block at rest pushed by a 280N horizontal force up to 30degree inclined planw with a roughness coefficient of 0.15


1
Expert's answer
2021-12-02T10:04:32-0500


1) Determine the normal force:


"N=mg\cos\theta+F\sin\theta."

2) The force of friction:


"f=\mu N=\mu(mg\cos\theta+F\sin\theta)."

3) The net force:


"F_\text{net}=F\cos\theta-f,\\\space\\
F_\text{net}=F\cos\theta-\mu(mg\cos\theta+F\sin\theta)=164\text{ N}."

4) Calculate the acceleration:

"a=F_\text{net}/m=3.64\text{ m/s}^2."